(1-w+w^2) (1-w^2+w^4) (1-w^4+w^8) to 2n factors is   
a)2 b)2^2n c)2n d)2^n

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
The ans woud be (x-1)^2n
w^1 = w 
w^2 = w
w^4 = w 
w^8 = w^6+w^2 = w^2
2^n = 3n , 3n + 1 , 3n +2
1+w+w^2 = 0 
=> (x-1)(x-1)(x-1)......2n 
=> (x-1)^2n is ans